OpenStudy (anonymous):

The diagonal of a square has a length of cm. Find the length of a side of the square.

OpenStudy (anonymous):

The diagonal of a square has a length of\[8\sqrt{2}\] cm. Find the length of a side of the square.

OpenStudy (khaakenajaf):

THE DIAGONAL OF A SQUARE=8√2 THUS YOU HAVE A RIGHT TRIANGLE WITH 2 EQUAL SIDES & A HYPOTENUSE OF 8 X^2+X^2=(8√2)^2 2x^2=64*2 x^2=64 x=64 cm

OpenStudy (khaakenajaf):

and this is the length

OpenStudy (anonymous):

A. \[2\sqrt{2}\] B. 4cm C. \[5\sqrt{2}\] D. 8cm

OpenStudy (anonymous):

@KhaakeNajaf

OpenStudy (anonymous):

divide by \(\sqrt2\) and get \(8\)